ECTDL SPECTRUM OF ^{15}NH_{3} IN THE 6370-6580 CM^{−1} REGION OF THE ν_{1} + ν_{3} NH-STRETCHING COMBINATION BAND

\underline{R.M. LEES}, Department of Physical Sciences, University of New Brunswick, Saint John, NB, Canada E2L 4L5.

The infrared spectrum of the N-15 isotopic species of ammonia is being investigated in the 6370-6580 cm^{−1} region with an external cavity tunable diode laser spectrometer. The patterns of some of the stronger absorption lines are very similar to those known for the ν_{1} + ν_{3} N-H stretching combination band of normal ^{14}NH_{3}, permitting immediate assignment by analogy with the known N-14 transitions. For ^{15}NH_{3}, the ν_{1} + ν_{3} band origin is shifted downwards by 12.2 cm^{−1} relative to ^{14}NH_{3}. From the preliminary spectra, the absorption intensities appear fairly comparable for both species, but more careful measurements to better define the isotopic intensity ratio are underway.